If one charge attracts the other with a force of 0.138N, what are the magnitudes of the two charges if their total charge is also 18.8 ?C? The charges are at a distance of 1.30 m apart. Note that you may need to solve a quadratic equation to reach your answer.
Answer #1


Given that F = 0.138 N

Q1+Q2 = 18.8*10^-6 C..(1)

F = k*Q1*Q2/r^2 = 0.138

Q1*Q2 = 0.138*r^2/k = (0.138*1.3^2)/(9*10^9) = 25.9*10^-12

Q1-Q2 = sqrt((Q1+Q2)^2-4*(Q1*Q2)) = sqrt((18.8*10^-6)^2-(4*25.9*10^-12)) = 15.8*10^-6.........(2)

solving (1) and(2) we get

2*Q1 = (18.8+15.8)*10^-6

Q1 = 17.3 uC

Q2 = 1.5 uC
Answer #2

Q1+Q2 = 18.8e-6 C


Q2 = 18.8e-6 - Q1 .............(1)


k*Q1*Q2/r^2 = F


(9e9*Q1*Q2)/(1.3*1.3) = 0.138

Q1*Q2 = 2.59e-11 ........(2)

substitutin 1 in 2


Q1*(18.8e-6-Q1) = 2.59e-11

Q1^2 - 18.8*10^-6Q1 + 2.59*10^-11 = 0

Q1 = 1.49*10^-6 C   <-----answer


Q2 = 18.8*10^-6 - 1.49*10^-6

Q2 = 17.31*10^-6 C        <--answer
Answer #4

Given,

Force of attraction of two charges = 0.138N

Total chare = 18.8 C and distance = r = 1.30m

Coulomb